OpenStudy (adamk):

What is cos(2x) if sin theta = 4/5 (0 < theta < pi/2). I got cos theta to be 5/sqrt(41) and sin2theta to be 8/sqrt(41) but I don't know how to do this one.

OpenStudy (solomonzelman):

You are given that `sin(x)=4/5` [with \(0<\theta<\pi/2\)], and that implies that adjacent side is 3. (Why? The hypotenuse is 5, and opposite side is 4, so... 5^2-4^2=9=3^2)

OpenStudy (solomonzelman):

So, your cosine should be 3/5.

OpenStudy (solomonzelman):

And for \(\cos(2\theta )\) use the following property: \(\cos(2\theta )=1-2\sin^2\theta\)

OpenStudy (solomonzelman):

(Showing, that all you needed is that property, without even having to find the cosine of theta .... )

OpenStudy (adamk):

@SolomonZelman My bad. For some reason I was thinking of sine as opposite divided by adjacent... aka tangent. Anyway the part I was stuck on was the 2sin^(2)theta.

OpenStudy (solomonzelman):

How to prove the property, or how to plug it in?

OpenStudy (adamk):

Plug it in

OpenStudy (solomonzelman):

\(\cos(2\theta )=1-2(\sin \theta)^2\quad \Longrightarrow \quad 1-2(4/5)^2=1-2(16/25)=1-32/25=-7/25\)

OpenStudy (solomonzelman):

the last part should be .... \(\color{blue}{-7/25}\)

OpenStudy (adamk):

Oh ok that easy. Thanks